OpenStudy (anonymous):

Solve. |2x + 4| - 5 = 6

OpenStudy (lgbasallote):

solve for x when 2x + 4 - 5 = 6 AND 2x + 4 - 5 = -6

OpenStudy (anonymous):

add 5 to both sides and solve \[|2x+4|=11\] \[2x+4=11\] or \[2x+4=-11\] etc

OpenStudy (anonymous):

careful here have to add 5 to both sides first

OpenStudy (lgbasallote):

oh...it's not the same? i assumed they were :/

OpenStudy (anonymous):

@lgbasallote try it and see. first one will not make a difference but there is a difference between \[2x + 4 - 5 = -6\] \[2x-1=-6\] \[2x=-5\] \[x=-\frac{5}{2}\] and \[2x+4=-11\] \[2x=-15\] \[x=-\frac{15}{2}\]

OpenStudy (lgbasallote):

uhmm yeah..i realized after i gave it some thought =))) ill remember that thanks